OpenStudy (anonymous):

x+3/3=y+2/2 then x/3=? anyone with a step by step solution?

OpenStudy (anonymous):

(x+3) / 3 = (y+2) / 2 (x/3) + (3/3) = (y+2) / 2 (x/3) + 1 = (y+2) /2 x/3 = [(y+2) / 2] - 1 x/3 = y/2

OpenStudy (anonymous):

o i see thanks again

OpenStudy (anonymous):

Welcome :)